# Beyond Numeric Awards: In-Context Dueling Bandits with LLM Agents

### v1

@@ 5.2 Theoretical Guarantees for IF-Enhanced LLM
For the dueling bandits problem with $K$ arms and time horizon $T$, there exists a preference structure and an attacker strategy with budget $\Phi(T)$, such that any standalone LLM agent, whose policy is represented by Eq.(1) and whose worst-case behavior under the original prompt satisfying Assumption 1, will suffer an expected regret of $\Omega\left(\min\left\{\Phi(T),\frac{T}{K}\right\}\right)$.
The proof of Theorem 5.1 is provided in Appendix A.2.2. The theorem underscores the suboptimality of standalone LLM agents in DB problems, particularly when input prompts are subjected to adversarial attacks. This vulnerability highlights the need for a more robust approach that harnesses the power of LLMs while offering theoretical guarantees under diverse prompting scenarios.
Expected Regret Bounds of IF-Enhanced LLM. Following the algorithm design of IF-Enhanced LLM introduced in Section 5.1, IF-Enhanced LLM inherits the theoretical guarantees of IF2 (see Appendix A.2.1), while leveraging the benefits of LLM-enhanced exploration. Specifically, IF-Enhanced LLM offers the following theoretical guarantee:
Theorem 5.2 (Expected Regret Bounds).
Suppose after $T_{\mathrm{LLM}}$ number of steps, the arms suggested by an LLM agent contain the best arm $b^{*}$. Under the same assumptions of IF2, with a hyper-parameter $\epsilon=\epsilon_{1,2}$, the expected strong regret of IF-Enhanced LLM satisfies that
$\mathbb{E}\left[\mathsf{SR}(\textsc{IF-Enhanced LLM})\right]\leq\min\left\{\widetilde{O}\left(T_{\mathrm{LLM}}+\frac{K\log K}{\epsilon_{1,2}}\right),\widetilde{O}\left(\frac{K\log K}{\epsilon_{1,2}}\log T\right)\right\},$
and the expected weak regret of IF-Enhanced LLM is bounded from above by
$\mathbb{E}\left[\mathsf{WR}(\textsc{IF-Enhanced LLM})\right]\leq\min\left\{O(T_{\mathrm{LLM}}),\widetilde{O}\left(\frac{K\log K}{\epsilon_{1,2}}\log T\right)\right\},$
@@
The proof of Theorem 5.2 is provided in Appendix A.2.2. The required assumptions are precisely stated in Appendix A.2.1. Theorem 5.2 establishes the efficacy and robustness of IF-Enhanced LLM:
Efficacy: Both the strong and weak regret for IF-Enhanced LLM are guaranteed to satisfy a worst-case upper bound of $\widetilde{O}\left(({K\log K}/{\epsilon_{1,2}})\log T\right)$, which is only worse than the well-known lower bound of $\smash{\Omega\left(({K}/{\epsilon_{1,2}})\log T\right)}$ in [3] by a factor of $\log K$. Additionally, as illustrated in Figures 28, and 11, IF-Enhanced LLM has the potential to identify the best arm after a short exploration stage. This results in strong and weak regret bounds of $\smash{\widetilde{O}(T_{\mathrm{LLM}}+({K}/{\epsilon_{1,2}})\log K)}$ and $O(T_{\mathrm{LLM}})$, respectively, provided the LLM agent suggests a pair of arms that includes the best arm $b^{*}$.
Robustness: The worst-case upper bounds on the strong and weak regret hold regardless of the specific prompting scenario, ensuring that IF-Enhanced LLM maintains its theoretical guarantees even in the presence of noisy or adversarial prompts, as considered in Theorem 5.1. This robustness is particularly important in practical applications, where the prompts provided to the LLM agent may be subject to various uncertainties or manipulations. By leveraging the strengths of LLMs while incorporating a robust algorithmic framework, IF-Enhanced LLM offers a reliable and theoretically grounded solution for DB problems in real-world settings, where the prompting quality may not always be optimal.
5.3 Empirical Evaluation of IF-Enhanced LLM
Regarding [Q2], we design a two-fold evaluation to assess efficacy and robustness. The evaluation is conducted on the Easy instance, which provides higher distinguishability, allowing us to observe convergence and regret differences within a practical number of steps. First, we compare the strong regret of IF-Enhanced LLM against state-of-the-art baseline algorithms to validate its efficacy. Second, we investigate the robustness of IF-Enhanced LLM with noisy and adversarial prompts.
@@ 5.3.1 Efficacy Evaluation: Strong Regret and Weak Regret
Figure 4: Comparisons between IF-Enhanced LLM, GPT-4 Turbo, IF2, and state-of-the-art baseline algorithms (Self-Sparring and DTS). Left and Right: strong and weak regret on the Easy instance.
Hyper-parameters. In our implementation of IF-Enhanced LLM (see Algorithm 1), there are two hyper-parameters: the threshold parameter $t$, which controls the maximum number of comparisons between arms, and the confidence parameter $\delta$, which determines the confidence level for pruning suboptimal arms. For the threshold parameter $t$, we considered values from the set $\{50,100,200\}$, and for the confidence parameter $\delta$, we explored values from $\{0.1,0.2,0.4\}$. After fine-tuning, we found that setting $t=50$ and $\delta=0.4$ provided the best performance in terms of cumulative regret. These hyper-parameter values strike a balance between the number of comparisons required to identify the best arm and the confidence level for pruning suboptimal arms, enabling IF-Enhanced LLM to efficiently explore and exploit the available arms in the dueling bandits setting.
We evaluate the cumulative strong and weak regret performance of the proposed IF-Enhanced LLM with different confidence parameter settings ($\delta=0.1,0.2,0.4$) and $t=50$:
Strong Regret. Figure 4 (Left) demonstrates that IF-Enhanced LLM exhibits competitive performance across different $\delta$ values. When $\delta=0.1$, IF-Enhanced LLM has slightly higher strong regret compared to the state-of-the-art baselines due to the conservative choice of $\delta$. As $\delta$ increases to 0.2 and 0.4, the cumulative expected strong regret of IF-Enhanced LLM decreases, eventually outperforming both state-of-the-art baselines when $\delta=0.4$. These results highlight a trade-off between robustness and optimality by selecting different confidence intervals in IF-Enhanced LLM. A smaller $\delta$ value demands higher confidence in decision-making, leading to more conservative exploration and potentially higher regret. Conversely, a larger $\delta$ allows for more aggressive exploration, enabling IF-Enhanced LLM to identify the optimal arm sooner, leading to reduced cumulative strong regret.
Weak Regret. Figure 4 (Right) showcases the superior weak regret performance of IF-Enhanced LLM. When $\delta=0.2$ and $\delta=0.4$, IF-Enhanced LLM consistently achieves lower cumulative weak regret, demonstrating its ability to effectively identify and include the optimal arm in its comparisons. As emphasized in Table 1, this can be attributed to GPT-4 Turbo’s ability to consistently identify the best arm.
@@ 5.3.2 Robustness Evaluation: Noisy and Adversarial Prompts
Figure 5: Robustness evaluation under prompt perturbations. The specific noisy (biased history) and adversarial (reversed goal) prompts are shown in Figure 7.
Recent studies [2, 37] have emphasized the importance of varying prompts to elicit the desired behavior from LLMs in decision-making tasks, highlighting the potential limitations of prompt quality. Results obtained from a single prompt template may lead to unreliable conclusions that cannot generalize to real-world situations where optimal prompts are often unavailable. Thus, we evaluate the robustness of IF-Enhanced LLM in DB tasks by employing two types of prompt perturbations (see Figure 7) along with the original prompt (see Figure 6).
Original Prompt. Under the initial prompt, IF-Enhanced LLM leverages the LLM’s ability to quickly identify the best arm through exploration. As shown in Figures 14 and 15 (Top Row), we observe that: (i) Compared to IF2, IF-Enhanced LLM benefits from the LLM’s exploration ability by initializing with the best arm as the incumbent when entering the IF2 phase. (ii) Compared to GPT-4 Turbo, convergence to the Condorcet winner is guaranteed for IF-Enhanced LLM with high probability.
Biased History. We inject an incorrect history into the prompt, where each non-optimal arm initially wins against the best arm 10 times, while keeping the underlying preference matrix unchanged. In this scenario, standalone LLM agents are observed to get trapped in local optima for extended periods. In contrast, IF-Enhanced LLM overcomes this limitation by employing uniform comparisons in the IF2 phase to escape such suboptimal exploration modes (see Figure 15 (Middle Row)). This effectively addresses the local optima problem highlighted in the disadvantages of standalone LLMs in Table 1.
Reversed Goal. When the prompt is adversarially modified from maximizing reward to minimizing, the LLM consistently recommends non-optimal arms after its exploration stage. Nevertheless, even with adversarial prompts, IF-Enhanced LLM still maintains a robust behavior and achieves near-optimal cumulative strong regret, as shown in Figure 15 (Bottom Row). Since the LLM’s exploration capability is only utilized within the bounded length of the Match Arms procedure, the impact of the reversed goal on the exploitation phase is eliminated.
Figure 5 presents the cumulative strong regret results comparing IF-Enhanced LLM against standalone LLM agents and the IF2 algorithm across three prompt designs. Notably, IF-Enhanced LLM with $\delta=1/(TK^{2})$ (consistent with IF2 to showcase the robust behavior) achieves near-optimal cumulative regret with low variance even with noisy and adversarial prompts, validating the regret bounds stated in Theorem 5.2. The experiment in this section is extended from 2000 steps to 3000 steps to illustrate GPT-4 Turbo’s non-convergence and growing regret. IF-Enhanced LLM and IF2 converge to the best arm within 2000 steps, while GPT-4 Turbo’s cumulative expected regret continues to increase, indicating the instability of standalone LLM agents.
@@ 6 Conclusion
This work evaluates LLM agents as in-context decision-makers for the problem of dueling bandits (DB) through case studies. We find that the top-performing LLM (GPT-4 Turbo) exhibits strong capabilities in identifying the Condorcet winner and achieving low weak regret while maintaining minimal variance. Yet, stemming from the fact that LLMs are primarily designed and trained for word token prediction rather than optimizing decision-making objectives, GPT-4 Turbo struggles to consistently converge to a single optimal solution in the long run and is vulnerable to prompt design variations. To address these limitations, we propose IF-Enhanced LLM, an LLM-augmented approach that takes advantages of both LLMs and classic DB algorithms, providing theoretical guarantees and robust empirical performance even with noisy and adversarial prompts. Our findings shed light on the capabilities and challenges of LLMs as in-context decision-makers, and pave the way for the employment of LLMs in complex decision-making tasks.
Limitations and Future Works. Due to computational constraints and financial costs, our case studies focus on a relatively small number of arms ($K=5$). To address this limitation, it is worth investigating the scalability of the proposed approach to larger decision spaces with more arms as certified by Theorem 5.2. In addition, while IF2 is employed as a representative Explore-Then-Exploit [18] algorithm to design an LLM-augmented algorithm in this work, utilizing other ongoing regret-minimization algorithms, such as RUCB [18], would be an interesting future direction. Moreover, considering more complex preference structures, such as non-transitive preferences, would provide insights on algorithm design. Investigating other variants of dueling bandits, including contextual dueling bandits, multi-dueling bandits, and adversarial dueling bandits, would further expand the applicability of this work. Finally, we have only experimented with some of the models available to us as of May 2024, selecting high-profile closed-source models over open-source models. Exploring the performance of a wider range of LLMs, including open-source models, would provide a more comprehensive understanding of the capabilities and limitations of LLMs in the dueling bandits setting.
Broader Impacts. The ability of LLMs to encode complex contextual information enables their application in various decision-making scenarios beyond the scope of traditional DB algorithms. This could lead to more informed and context-aware decision-making in domains such as healthcare, finance, and recommender systems, where incorporating user feedback and adaptivity is crucial for handling dynamic relative preferences. However, the substantial computational resources required to train and deploy large-scale LLMs for decision-making tasks raise concerns about energy consumption and environmental impact, especially when dealing with a large set of arms.
@@
This appendix provides supplementary information and additional experimental results to support the main text. The content is organized into two main parts:
  • Theoretical Part: Algorithm Design and Analysis of IF-Enhanced LLM
    • Appendix A.1 describes the IF-Enhanced LLM algorithm stated in Section 5.1, detailing its key features and implementation remarks.
    • Appendix A.2.1 presents the necessary assumptions and lemmas for the theoretical analysis of IF-Enhanced LLM in Section 5.2.
    • Appendix A.2.2 proves Theorem 5.1 and 5.2, establishing IF-Enhanced LLM’s regret bounds.
@@
    • Appendix B.1.1 illustrates the prompt design and prompt perturbations logic.
    • Appendix B.1.2 provides a five consecutive time steps example of GPT-4 Turbo to showcase its behavior.
@@
    • Appendix B.2 presents supplementary experimental results, providing further insights into the performance and behavior of the algorithms in Sections 4 and 5.
Appendix A Algorithm Design and Analysis of IF-Enhanced LLM

### v3

@@ 4.2 Theoretical Guarantees for LEAD
For the dueling bandits problem with $K$ arms and time horizon $T$, there exists a preference structure and an attacker strategy with budget $\Phi(T)$, such that any standalone LLM agent, whose policy is represented by Eq.(1) and whose worst-case behavior under the original prompt satisfying Assumption 4, will suffer an expected regret of $\Omega\left(\min\left\{\Phi(T),{T}/{K}\right\}\right)$.
The proof of Theorem 4.1 is provided in Appendix B.3.2. The theorem underscores the suboptimality of standalone LLM agents in DB problems, particularly when input prompts are subjected to adversarial attacks. This vulnerability highlights the need for a more robust approach to use in-context LLM agents while offering theoretical guarantees under diverse prompting scenarios.
Regret Bounds. Following the algorithmic design of LEAD in Section 4.1, LEAD (IF2 base) inherits the theoretical guarantees of IF2 (see Appendix B.3.1), while nontrivially leveraging the benefits of LLMs’ exceptional weak regret performance for exploration across a range of instances within RDB. Specifically, LEAD (IF2 base) has the following theoretical guarantee:
Theorem 4.2 (Expected Regret).
Suppose for $t\geq T_{\mathrm{LLM}}$, the arms recommended by an LLM agent contain the best arm $b^{*}$. Under Assumptions 1-3, the expected strong regret of LEAD (IF2 base) satisfies $\mathbb{E}\left[\mathsf{SR}(\textsc{LEAD})\right]\leq\widetilde{O}\left({(K\log T)}/{\epsilon_{1,2}}\right),$ and the expected weak regret can be bounded by
$\mathbb{E}\left[\mathsf{WR}(\textsc{LEAD})\right]\leq\min\left\{\widetilde{O}\left(T_{\mathrm{LLM}}+\frac{K\log K}{\epsilon_{1,2}}\right),\widetilde{O}\left(\frac{K\log T}{\epsilon_{1,2}}\right)\right\},$ (3)
@@
Note that Theorem 4.2 is general such that we do not assume any specific adversarial behaviors of the LLM agent, including Assumption 4. The proof of Theorem 4.2 is provided in Appendix B.3.2. The required assumptions are precisely stated in Appendix B.3.1. Theorem 4.2 establishes a best-of-both-worlds result in terms of the efficacy and robustness of LEAD.
Efficacy. As illustrated in Figures 23, and  9, LEAD has the potential to identify the best arm after a short exploration stage. This results in strong and weak regret bounds of $\smash{\widetilde{O}(T_{\mathrm{LLM}}+({K}/{\epsilon_{1,2}})\log K)}$ and $O(T_{\mathrm{LLM}})$, respectively, that are independent of the horizon length $T$, provided the LLM agent suggests a pair of arms that includes the best arm $b^{*}$. Furthermore, when the prompt contains extra textual context that can infer the relative preferences between arms, $T_{\mathrm{LLM}}$ will become smaller, further enhancing the best-case performance. We consider it an important direction for future work within the Contextual Dueling Bandit framework Dudík et al. (2015).
Guaranteed Convergence. Additionally, both the strong and weak regret for LEAD are guaranteed to satisfy a worst-case upper bound of $\smash{\widetilde{O}\left(({K}/{\epsilon_{1,2}})\log T\right)}$, which is only worse than the information-theoretic lower bound of $\smash{\Omega\left(({K}/{\epsilon_{1,2}})\log T\right)}$ in (Yue et al., 2012) by a poly-logarithmic factor of $T$. The worst-case upper bounds on the strong and weak regret hold regardless of the specific prompting scenario, ensuring that LEAD maintains its theoretical guarantees even in the presence of noisy or adversarial prompts, as considered in Theorem 4.1. This safety guarantee is particularly important in practical applications, where the prompts provided to the LLM agent may not always be optimal.
The following theorem indicates that the additional term $({K\log K})/{\epsilon_{1,2}}$ in (3) is almost tight. Its proof is provided in Appendix B.3.2.
Theorem 4.3 (Converse).
Given any algorithm ALG for dueling bandits provided with an in-context LLM agent recommending arms, if it satisfies $\mathbb{E}\left[\mathsf{WR}(\textsc{ALG})\right]\leq T_{\mathrm{LLM}}$ for all $T_{\mathrm{LLM}}$, then it must hold ${\mathbb{E}\left[\mathsf{SR}(\textsc{ALG})\right]\geq\mathbb{E}\left[\mathsf{WR}(\textsc{ALG})\right]\geq\Omega\left(T\right)}$ for some instance of the LLM agent.
4.3 Empirical Evaluation of LEAD
Regarding [Q2], we design a two-fold evaluation to assess efficacy and robustness. The evaluation is conducted on the Transitive-Easy instance, which provides higher distinguishability, allowing us to observe convergence and regret differences within a practical number of steps. First, we compare the strong and weak regret of LEAD against state-of-the-art baseline algorithms to validate its efficacy. Second, we investigate the robustness of LEAD with noisy and adversarial prompts.
@@ 4.3.1 Efficacy Evaluation: Strong Regret and Weak Regret
Figure 5: Comparisons between LEAD, GPT-4 Turbo, and baseline algorithms (IF2, Self-Sparring and DTS). Left and Middle: strong and weak regret on the Transitive-Easy instance. Right: robustness evaluation under prompt perturbations (prompts are in Appendix C.1.2).
Hyper-parameters. In our implementation of LEAD (see Algorithm 1), there are two hyper-parameters: the threshold parameter $t$, which controls the maximum number of comparisons between arms, and the confidence parameter $\delta$, which determines the confidence level for pruning suboptimal arms. For the threshold parameter $t$, we considered values from the set $\{50,100,200\}$, and for the confidence parameter $\delta$, we explored values from $\{0.1,0.2,0.4\}$. After fine-tuning, we found that setting $t=50$ and $\delta=0.4$ provided the best performance in terms of cumulative regret.
We evaluate the cumulative strong and weak regret performance of the proposed LEAD with different confidence parameter settings ($\delta=0.1,0.2,0.4$) and $t=50$: Figure 5 (Left and Middle) demonstrates that LEAD exhibits competitive performance across different $\delta$ values. For strong regret, $\delta=0.1$ results in more conservative exploration, leading to slightly higher regret compared to baselines. As $\delta$ increases ($\delta=0.2$ or $0.4$), LEAD achieves lower cumulative strong regret, outperforming all the baselines at $\delta=0.4$ due to more aggressive exploration to identify the optimal arm sooner. Similarly, for weak regret, LEAD consistently achieves superior performance. When $\delta=0.2$ and $\delta=0.4$, LEAD effectively identifies and includes the optimal arm in comparisons. These hyper-parameter values strike a balance between the number of comparisons required to identify the best arm and the confidence level for pruning suboptimal arms, enabling LEAD to efficiently explore and exploit the available arms in-context for the dueling bandits setting.
@@ 4.3.2 Robustness Evaluation: Noisy and Adversarial Prompts
Recent studies (Loya et al., 2023; Krishnamurthy et al., 2024) have emphasized the importance of varying prompts to elicit the desired behavior from LLMs in decision-making tasks, highlighting the potential limitations of prompt quality. Results obtained from a single prompt template may lead to unreliable conclusions that cannot generalize to real-world situations where optimal prompts are often unavailable. Thus, we evaluate the robustness of LEAD by employing two types of prompt perturbations (see Figure 8) along with the original prompt (see Figure 7). Across all scenarios, LEAD demonstrates superior performance and robustness compared to standalone GPT-4 Turbo.
Figure 6: Duel selection trajectory of GPT-4 Turbo and LEAD under different settings (Figures 7 and 8). Top: original prompt. Middle: noisy prompt (biased history). Bottom: adversarial prompt (reversed goal).
Original Prompt. Under the initial prompt, LEAD leverages the LLM’s ability to quickly identify the best arm through exploration (under the DB instance within RDB Eq.(2)). As shown in Figure 6 (Top Row), we observe that LEAD benefits from the LLM’s exploration ability by initializing with the best arm as the incumbent when entering the IF2 phase. Compared to GPT-4 Turbo, convergence to the Condorcet winner is guaranteed for LEAD with high probability.
Biased History. We inject an incorrect history into the prompt, where each non-optimal arm initially wins against the best arm 10 times, while keeping the underlying preference matrix unchanged. LLM agents are observed to get trapped in local optima for extended periods, where LEAD overcomes this limitation by employing uniform comparisons in the IF2 phase to escape such suboptimal exploration modes.
Reversed Goal. When the prompt is adversarially modified from maximizing reward to minimizing, the LLM consistently recommends non-optimal arms after its exploration stage. Even with adversarial prompts, LEAD still achieves near-optimal cumulative strong regret. Since the LLM’s exploration capability is only utilized within the bounded length of the Match Arms procedure, the impact of the reversed goal on the exploitation phase is mitigated.
Figure 5 (right) presents the cumulative strong regret results comparing LEAD against standalone LLM agents and the IF2 algorithm across three prompt designs. Notably, LEAD with $\delta=1/(TK^{2})$ (consistent with IF2 to showcase the robust behavior) achieves near-optimal cumulative regret with low variance even with noisy and adversarial prompts, validating the regret bounds stated in Theorem 4.2. LEAD and IF2 converge to the best arm within 2000 steps, while GPT-4 Turbo’s cumulative expected regret continues to increase, indicating the instability of standalone in-context LLM agents.
@@ 5 Conclusion
This paper evaluates LLMs as in-context decision-makers for standard context-free dueling bandits (DB) with a Condorcet Winner, offering the first systematic insights into their strengths and limitations. Our findings reveal that LLMs’ decision-making in DB, driven by linguistic priors, achieves exceptional weak regret performance across both transitive and intransitive environment instances in the short-term. However, LLMs lack the necessary criteria for convergence and long-term performance generalization to hard scenarios, leading to an optimality gap between LLMs and classic DB algorithms in terms of strong regret. To bridge this gap, we propose LEAD, an agentic flow framework that integrates off-the-shelf DB algorithms with LLM agents through fine-grained adaptive interplay. This framework provides theoretical guarantees and demonstrates robust performance even under noisy and adversarial prompts. Our work contributes to the problem of In-Context Reinforcement Learning (ICRL). The framework we proposed sheds light on how language-based reasoning can inspire robust frameworks that translate words into actions, paving the way for more trustworthy AI systems through the interplay between rule-based experts and in-context LLM agents.
@@
This appendix provides supplementary information and additional experimental results to support the main text. The content is organized into three main parts:
  1. A.
    Related Works
  2. B.
    Theoretical Part: Algorithm Design and Analysis of LEAD
    • Appendix B.1 presents the algorithm design logic using Explore-then-Exploit methods.
    • Appendix B.2 describes the LEAD algorithm stated in Section 4.1, detailing its key features and implementation remarks.
    • Appendix B.3.1 presents the necessary definitions, assumptions and lemmas for the theoretical analysis of LEAD in Section 4.2.
    • Appendix B.3.2 proves Theorem 4.14.2, and 4.3, establishing LEAD’s regret bounds.
  3. C.
@@
    • Appendix C.1.1 illustrates the transitive and intransitive environments construction.
    • Appendix C.1.2 illustrates the prompt design and prompt perturbations logic.
    • Appendix C.1.3 provides exemplars of GPT-4 Turbo to showcase their behavior.
@@
    • Appendix C.2 presents supplementary experimental results, providing further insights into the performance and behavior of the algorithms in Sections 3 and 4.
Appendix A Related Works
We provide the detailed related works as follows.
Dueling Bandits. The problem of dueling bandits was initially introduced in [Yue et al., 2012]. Various methods have been proposed to tackle the task since then. These methods can be broadly classified into two categories as Explore-Then-Exploit methods and Ongoing Regret Minimization methods according to [Zoghi et al., 2014a]. Explore-Then-Exploit methods focus on identifying the best arm with high confidence before exploiting it, such as Interleaved Filter (IF[Yue et al., 2012] and Beat the Mean (BTM[Yue and Joachims, 2011], etc. In contrast, Ongoing Regret Minimization methods explicitly target the objective of minimizing cumulative regret, including Relative Upper Confidence Bound (RUCB[Zoghi et al., 2014b] and Self-Sparring [Sui et al., 2017], etc. Dueling bandit problem and preference feedback in general has a wide variety of applications, including recommendation systems [Yue et al., 2012], robotics [Tucker et al., 2020], and most recently, the training algorithm of large language models, such as Reinforcement Learning from Human Feedback (RLHF) [Ouyang et al., 2022].
LLM Agents for Multi-Armed Bandits. Several recent works have explored evaluating the capabilities of LLMs in bandit problems. For example, [Baheri and Alm, 2023] proposed an approach to enhance contextual bandits by integrating LLMs as encoders. The LLMs’ ability to capture rich semantic and syntactic information from textual contexts is leveraged to provide the algorithm with a more informative representation of the context. The LLM-augmented algorithm transforms the raw context into a latent space vector using the LLM’s encoding capabilities. This encoded context is then used to guide the decision-making process. [Krishnamurthy et al., 2024] investigates whether LLMs can engage in exploration in simple MAB environments without additional training. They compared various prompt designs and found that GPT-4 with zero-shot chain-of-thought (CoT) reasoning and an externally summarized interaction history performed the best, while other configurations failed in exploration, either by never selecting the best arm after initial rounds or by selecting all arms nearly equally often. Different from the previous results, in this work we go beyond the settings of numeric rewards and investigate the capabilities of LLMs under preference feedback.
In-Context LLMs for Decision-Making. Beyond bandit problems, LLM agents have demonstrated strong capabilities in complex reasoning across a wide range of in-context reinforcement learning and decision-making tasks [Laskin et al., 2022, Lee et al., 2024, Zhou et al., 2022, Yao et al., 2024]. Various existing works aim to understand LLM agents’ capabilities for in-context decision-making, with notable examples including planning [Huang et al., 2022, Hao et al., 2023]. Additionally, LLM agents have been shown to enhance embodied agents in robotic applications by providing advanced task planning abilities [Brohan et al., 2023] and reward designing [Ma et al., 2023], further enabling the development of lifelong learning agents [Wang et al., 2023]. Besides these empirical successes, the authors of [Park et al., 2024] analyzed LLMs’ interactions in online learning and game theory settings through the lens of the regret metrics. They identified simple cases where LLMs fail to be no-regret. Another line of research incorporates LLMs into classic decision-making frameworks to create LLM-augmented online decision-makers. For instance, Liu et al. [Liu et al., 2024] utilized LLMs to enhance the components of warm starting, sampling candidates, and surrogate modeling in Bayesian optimization. Our work contributes to this broad area by integrating LLM agents with the classic Explore-then-Exploit DB algorithms to enhance the utilization of preference feedback.
Appendix B Algorithm Design and Analysis of LEAD